OpenStudy (anonymous):

NEED QUICK HELP PLEASE. WILL GIVE A MEDAL!! Francisco and Ryan are stuck simplifying radical expressions. Francisco has simplified to the quantity of x to the one half power, over x to the three eighteenth power. Ryan has simplified to the twenty seventh root of the quantity of x to the second times x to the third times x to the fourth. Using full sentences, describe how to fully simplify Francisco and Ryan's expressions. Describe if Francisco and Ryan started with equivalent expressions or if they started with expressions that are not equal.

OpenStudy (misty1212):

hi!

OpenStudy (misty1212):

if you use the equation tool, it will be readable

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

all u have to do is to subtract the exponents..

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is this two different questions?

OpenStudy (anonymous):

It is one question and how do i do that @magepker728

OpenStudy (misty1212):

what is \[\huge \frac{1}{2}-\frac{3}{18}\]

OpenStudy (misty1212):

btw just for the record \(\frac{3}{18}=\frac{1}{6}\) so it is really \[\frac{1}{2}-\frac{1}{6}\]

OpenStudy (anonymous):

would you take it and make it 1/2 + 6/1 then?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

just subtract like you always do with fractions it is a subtraction, not a division

OpenStudy (anonymous):

1/4 ?

OpenStudy (misty1212):

not sure how you got that number did you guess?

OpenStudy (anonymous):

no I tried to subtract

OpenStudy (misty1212):

do you know how to add and subtract fractions? like finding a common denominator and all that?

OpenStudy (anonymous):